# The canonical lamination calibrated by a cohomology class

### v3

@@ Abstract.
Let $\rho$ be a unit cohomology class of degree $d-1$, on a closed oriented Riemannian manifold of dimension $d$. We construct a lamination $\lambda_{\rho}$ whose leaves are exactly the minimal hypersurfaces calibrated by every calibration in $\rho$. The geometry of $\lambda_{\rho}$ is closely related to the geometry of the unit ball of $H_{d-1}(M,\mathbb{R})$ when it is equipped with Gromov’s stable norm, so our main theorem constrains the shape of the stable unit ball in terms of the topology of $M$. These results establish a close analogy between the stable norm and Thurston’s earthquake norm on the tangent space to Teichmüller space.
@@ 1. Introduction
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The stable norm $\|\alpha\|_{1}$ of a homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ is the infimum of the area of all $d-1$-cycles representing $\alpha$. The stable norm was introduced by Federer in his work [Federer1974], on the duality between area-minimizing currents and calibration cochains. Among other applications, the stable norm was studied by Gromov, [gromov2007metric], for its connections to systolic geometry and Brock and Dunfield, [Brock2017], because of its connection to the Thurston–Gromov simplicial norm.
A quarter-century ago, Auer and Bangert released a research announcement [Auer01], which proposed to study codimension-$1$ measured oriented laminations $\lambda$ in $M$ which minimize their mass in their homology class $[\lambda]\in H_{d-1}(M,\mathbf{R})$. [^1] We review the basic definitions related to laminations in §2.4. In our convention, all laminations are Lipschitz and nonempty. [^2] Some of Auer and Bangert’s work appears in an unpublished manuscript, [Auer12]. In codimension $1$, every homology class $\alpha$ can be represented by a mass-minimizing lamination $\lambda$ (whose mass then equals $\|\alpha\|_{1}$), which one can think of roughly think of as a canonical choice of representative of $\alpha$. If two laminations have common leaves, those leaves cannot intersect, and Auer and Bangert proposed to use this observation to establish a deep connection between the intersection theory of $M$ and the geometry of the stable unit ball. A similar approach was used by Balacheff and Massart, [Massart96, Massart2007], to study the stable unit ball when $M$ is a negatively curved surface.
While trying to prove the theorems claimed in [Auer01], it is often useful to imitate ideas of the works [Thurston98, Wolpert82, Gu_ritaud_2017, huang2024earthquakemetricteichmullerspace] of the Thurston school on Thurston’s asymmetric metric on Teichmüller space. To make this precise, let $g\geq 2$, let $\Sigma_{g}$ be the closed oriented surface of genus $g$, let $\mathscr{T}_{g}$ be its Teichmüller space, let $\rho,\sigma\in\mathscr{T}_{g}$ be hyperbolic metrics on $\Sigma_{g}$, and let $L(\rho,\sigma)\geq 1$ be the infimum of Lipschitz constants of maps $(\Sigma_{g},\rho)\to(\Sigma_{g},\sigma)$ homotopic to $\operatorname{id}_{\Sigma_{g}}$. Thurston’s stretch metric on $\mathscr{T}_{g}$ is $\log L$. Thurston’s stretch metric is studied using geodesic laminations on $(\Sigma_{g},\rho)$, and in particular the canonical maximally-stretched lamination given by the following theorem.
Theorem 1.1 ([Gu_ritaud_2017]).
@@
The costable norm, $\|\cdot\|_{\infty}$, on $H^{d-1}(M,\mathbf{R})$ is the dual norm of the stable norm. The purpose of this paper is to flesh out the idea that the costable norm of $M$ is closely analogous to Thurston’s stretch distance in an infinitesimal neighborhood of a hyperbolic metric. Our main theorem is that for each class of unit costable norm, there is a lamination in $M$ of minimal hypersurfaces which is analogous to Thurston’s canonical lamination. Studying the structure of this lamination yields several of the proposed theorems of [Auer01]. A sample application of the theory we shall develop is that under purely topological assumptions, $M$ has many uniquely ergodic laminations of minimal hypersurfaces.
A calibration (of codimension $1$) is a closed $d-1$-form $F$ such that $\|F\|_{L^{\infty}}=1$. For any calibration $F$ on $M$, a hypersurface $N\subset M$ is $F$-calibrated if the pullback of $F$ to $N$ is the area form on $N$. If $N$ is $F$-calibrated, then the mean curvature of $N$ is 00, and if $N$ is closed then $N$ minimizes its area in its real homology class. This brings us to our main theorem:
@@
The lamination $\lambda_{\rho}$ is the canonical lamination calibrated by $\rho$. We prove Theorem 1.2 in §4. The proof uses multiple results from [BackusCML], including the interpretation of mass-minimizing laminations in terms of functions of least gradient, functions $u$ which minimize their total variation $\int_{M}\star|\mathop{}\!\mathrm{d}u|$. We carefully note that the lamination $\lambda_{\rho}$ is not itself mass-minimizing, since it may not admit any sort of transverse measure and therefore does not have a well-defined homology class. However, any measured sublamination of $\lambda_{\rho}$ is mass-minimizing in its homology class.
To illustrate Theorem 1.2, suppose that $d=2$, so that we can identify homotopy classes of maps $M\to\mathbf{S}^{1}$ with homomorphisms $\pi_{1}(M)\to\mathbf{Z}$, which in turn can be identified with lattice points in $H^{1}(M,\mathbf{R})$. Let $\rho$ be such a lattice point; by rescaling $M$ we may assume that $\|\rho\|_{\infty}=1$. In that case, any calibration which represents $\rho$ is the derivative of a minimizing Lipschitz map in the homotopy class $\rho$, and every leaf of the canonical lamination calibrated by $\rho$ is maximally stretched by every minimizing Lipschitz map in $\rho$. Thus Theorem 1.2 is a generalization of a version of Theorem 1.1 where one works with homotopy classes $M\to\mathbf{S}^{1}$ rather than $[\operatorname{id}_{\Sigma_{g}}]$.
The definition of the costable norm makes sense in any cohomology group $H^{k}(M,\mathbf{R})$ and so it is natural to ask if Theorem 1.2 holds for $k\leq d-2$. If $k=1$, the analogue of Theorem 1.2 follows from Daskalopoulos and Uhlenbeck’s work [daskalopoulos2020transverse] on the $\infty$-Laplacian. However, if $22\leq k\leq d-$ 2, then the submanifolds calibrated by $\rho\in H^{k}(M,\mathbf{R})$ can intersect each other. For example, let $M=\mathbf{P}^{2}_{\mathbf{C}}$, let $\rho$ be the Kähler class of $\mathbf{P}^{2}_{\mathbf{C}}$, and let $F$ be the Kähler form of $\mathbf{P}^{2}_{\mathbf{C}}$, which is a calibration by Wirtinger’s inequality.
In §5, we study measured sublaminations of canonical calibrated laminations. This is motivated by the fact that the earthquake norm, the dual of the norm induced by Thurston’s stretch metric, is not strictly convex, and its failure to be strictly convex detects the failure of geodesic laminations to be uniquely ergodic [huang2024earthquakemetricteichmullerspace]. This suggests that if the stable norm is not strictly convex, then there should be canonical calibrated laminations which are not uniquely ergodic; it turns out that this is exactly what happens.
A finite Borel measure $\mu$ is transverse to a lamination $\lambda$, if $\operatorname{supp}\mu=\operatorname{supp}\lambda$ and $\mu$ is invariant under deformations which preserve the area forms of every leaf of $\lambda$. A transverse probability measure $\mu$ is ergodic if, for every Borel set $E$ which is a union of leaves of $\lambda$, either $\mu(E)=0$ or $\mu(E)=1$. The lamination $\lambda$ is uniquely ergodic, if there is a unique probability measure which is transverse to $\lambda$.
@@
Auer and Bangert [Auer01] observed that one can use a lemma of Arnoux and Levitt [Arnoux1986] to estimate the number of ergodic measures on sublaminations of a lamination without closed leaves. So by Corollary 1.3, the Arnoux–Levitt lemma applied to $\lambda_{\rho}$ determines the structure of $\rho^{*}$. A homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ has rational direction if there exists $c>0$ such that $c\alpha$ is in the image of the map $H_{d-1}(M,\mathbf{Z})\to H_{d-1}(M,\mathbf{R})$. Let $b_{1}(M):=\dim H_{1}(M,\mathbf{Q})$ be the first Betti number.
@@
For example, suppose that $M\cong\Sigma_{g}$ where $g\geq 2$. By a theorem of Massart [Massart1997StableNO], if a maximal flat $\rho^{*}$ has a point of rational direction, then $\rho^{*}$ has at most $33g-$ 3 vertices, all of which have rational direction. It follows that $\lambda_{\rho}$ consists of at most $33g-$ 3 closed geodesics, plus possibly a “spiraling” part which admits no transverse measures. On the other hand, if $\rho^{*}$ has no points of rational direction, then $\rho^{*}$ has at most $12g-$ 1 vertices, and $\lambda_{\rho}$ has no closed leaves.
@@
The analysis of $\lambda_{\rho}$ yields the following theorem on the strict convexity of the stable unit ball, which was proposed without proof by Auer and Bangert [Auer01, Theorems 6 and 7]. The intersection product $\alpha\cdot\beta$ of two homology classes $\alpha,\beta$ is the Poincaré dual of the cup product $\operatorname{PD}(\alpha)\smile\operatorname{PD}(\beta)$, and the derived series of a group $\Gamma$ is defined by letting $\Gamma^{(0)}:=\Gamma$ and $\Gamma^{(n+1)}$ be the commutator subgroup of $\Gamma^{(n)}$.
@@
    If there is a line segment $[\alpha,\beta]\subset\partial B$, then $\alpha\cdot\beta=0$.
@@
For example, suppose that $M$ has the homotopy type of a torus. Then $B$ is strictly convex and so $M$ has many uniquely ergodic laminations of minimal hypersurfaces. If $M=\mathbf{R}^{d}/[0,1]^{d}$, then these laminations are the irrational foliations of $M$, but what Theorem 1.6 says is that the same behavior occurs regardless of the Riemannian metric on $M$.
Since the canonical calibrated lamination $\lambda_{\rho}$ does not have to be uniquely ergodic, it is natural to ask if there is a canonical measure on $\lambda_{\rho}$. In upcoming work [daskalopoulos2025], Daskalopoulos and Uhlenbeck will show there is a favored measure on any canonical maximally stretched lamination with only closed leaves. The same proof works on $\lambda_{\rho}$, but it is quite lengthy, so we omit the proof. For each $p<\infty$, there is a unique representative $F_{p}\in L^{p}(M,\Omega^{d-1})$ of $\rho$ which is $p$-harmonic; that is,
@@
After taking a subsequence, $u_{p}$ converges in $L^{1}_{\mathrm{loc}}(\tilde{M})$ to a function $u$ of least gradient such that $\mu:=|\mathop{}\!\mathrm{d}u|\in\mathcal{M}(\lambda_{\rho})$. If $M$ is a hyperbolic surface and $\lambda_{\rho}$ only has closed leaves, then $\mu$ is independent of the choice of subsequence.
@@
Theorems 1.2 and 1.7 are explicitly based on theorems about the earthquake norm, and there are also versions of Theorem 1.4 and 1.6 for the earthquake norm proven in [huang2024earthquakemetricteichmullerspace]. In §5.6 we conjecture a version of Corollary 1.3 for the earthquake norm.
@@ Acknowledgments.
This work was closely inspired by ideas in [Auer01] of Franz Auer and Victor Bangert; I am especially grateful to Victor Bangert for allowing me to view their unpublished manuscript [Auer12]. I also thank Georgios Daskalopoulos and Karen Uhlenbeck for helpful discussions, and James Farre, Yi Huang, Zhenhua Liu, and Ben Lowe for helpful comments on an earlier draft. This research was supported by the National Science Foundation’s Graduate Research Fellowship Program under Grant No. DGE-2040433.
@@ 2.1. Notation
Unless otherwise noted, $M$ always denotes a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The operator $\star$ is the Hodge star on $M$. We denote the musical isomorphisms by $\sharp$,$\flat$ , ♭. We write $H^{\ell}$ for de Rham cohomology, but never a Sobolev space, which we instead denote $W^{\ell,p}$. The second fundamental form of a submanifold $N$ is $\mathrm{I\!I}_{N}$. If $K$ is a closed compact subset of a topological vector space, $\mathcal{E}(K)$ is the set of extreme points of $K$.
@@ 2.2. Differential forms in $L^{\infty}$
Let $x\in M$, and $10\leq k\leq d-$ 1. Then there exists $r_{*}>0$ which depends only on $\mathrm{Riem}_{M}$ near $x$ and the injectivity radius of $x$, such that for every $0<r\leq r_{*}$ and $F\in L^{\infty}(B(x,r),\Omega^{k+1}_{\rm cl})$, there exists a Hölder continuous $k$-form $A$ such that $F=\mathop{}\!\mathrm{d}A$.
@@
We may choose $r_{*}$ so that the exponential map $B_{\mathbf{R}^{d}}(0,r_{*})\to B(x,r_{*})$ is a diffeomorphism which induces topological isomorphisms for every function space under consideration. Thus it is no loss to replace $B(x,r)$ with the unit euclidean ball $\mathbf{B}^{d}$. By the main theorem of [Costabel2010], for every $1<p<\infty$ there is a continuous right inverse to the exterior derivative
@@
The next result is a rephrasing of [Anzellotti1983, Theorem 1.2], and asserts that closed $L^{\infty}$ $d-1$-forms can be integrated along Lipschitz hypersurfaces.
@@
Let $\iota:N\to M$ be the inclusion of an oriented Lipschitz hypersurface. Let $\mathcal{X}$ be the space of $F\in L^{\infty}(M,\Omega^{d-1})$ such that the components of $\mathop{}\!\mathrm{d}F$ are Radon measures. Then the pullback $\iota^{*}$ of $d-1$-forms extends to a bounded linear map
@@
where the supremum ranges over all oriented $k$-dimensional submanifolds $\Sigma$. It is clear that $\|F\|_{L^{\infty}_{*}}\leq\|F\|_{L^{\infty}}$, but if $F$ is a $d-1$-form, then the converse holds as well; we shall often use this fact without comment.

### v4

@@ Abstract.
Let $\rho$ be a unit cohomology class of degree d1d-1, on a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. We construct a lamination $\lambda_{\rho}$ whose leaves are exactly the minimal hypersurfaces calibrated by every calibration in $\rho$. The geometry of $\lambda_{\rho}$ is closely related to the geometry of the unit ball of $H_{d-1}(M,\mathbb{R})$ when it is equipped with Gromov’s stable norm, so our main theorem constrains the shape of the stable unit ball in terms of the topology of $M$. These results establish a close analogy between the stable norm and Thurston’s earthquake norm on the tangent space to Teichmüller space.
@@ 1. Introduction
Let $M$ be a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The stable norm $\|\alpha\|_{1}$ of a homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ is the infimum of the area of all d1d-1-cycles representing $\alpha$. The stable norm was introduced by Federer in his work [22], on the duality between area-minimizing currents and calibration cochains. Among other applications, the stable norm was studied by Gromov, [27], for its connections to systolic geometry and Brock and Dunfield, [11], because of its connection to the Thurston–Gromov simplicial norm.
A quarter-century ago, Auer and Bangert released a research announcement [3], which proposed to study codimension-$1$ measured oriented laminations $\lambda$ in $M$ which minimize their mass in their homology class $\lambda]\in H_{d-1}(M,\mathbf{R})$. [^1] We review the basic definitions related to laminations in §2.4. In our convention, all laminations are Lipschitz and nonempty. [^2] Some of Auer and Bangert’s work appears in an unpublished manuscript, [4]. In codimension $1$, every homology class $\alpha$ can be represented by a mass-minimizing lamination $\lambda$ (whose mass then equals $\|\alpha\|_{1}$), which one can think of roughly think of as a canonical choice of representative of $\alpha$. If two laminations have common leaves, those leaves cannot intersect, and Auer and Bangert proposed to use this observation to establish a deep connection between the intersection theory of $M$ and the geometry of the stable unit ball. A similar approach was used by Balacheff and Massart, [37, 7], to study the stable unit ball when $M$ is a negatively curved surface.
While trying to prove the theorems claimed in [3], it is often useful to imitate ideas of the works [46, 47, 28, 31] of the Thurston school on Thurston’s asymmetric metric on Teichmüller space. To make this precise, let $g\geq 2$, let $\Sigma_{g}$ be the closed oriented surface of genus $g$, let $\mathscr{T}_{g}$ be its Teichmüller space, let $\rho,\sigma\in\mathscr{T}_{g}$ be hyperbolic metrics on $\Sigma_{g}$, and let $L(\rho,\sigma)\geq 1$ be the infimum of Lipschitz constants of maps $(\Sigma_{g},\rho)\to(\Sigma_{g},\sigma)$ homotopic to $\operatorname{id}_{\Sigma_{g}}$. Thurston’s stretch metric on $\mathscr{T}_{g}$ is $\log L$. Thurston’s stretch metric is studied using geodesic laminations on $(\Sigma_{g},\rho)$, and in particular the canonical maximally-stretched lamination given by the following theorem.
Theorem 1.1 ([28]).
@@
The costable norm, $\|\cdot\|_{\infty}$, on $H^{d-1}(M,\mathbf{R})$ is the dual norm of the stable norm. The purpose of this paper is to flesh out the idea that the costable norm of $M$ is closely analogous to Thurston’s stretch distance in an infinitesimal neighborhood of a hyperbolic metric. Our main theorem is that for each class of unit costable norm, there is a lamination in $M$ of minimal hypersurfaces which is analogous to Thurston’s canonical lamination. Studying the structure of this lamination yields several of the proposed theorems of [3]. A sample application of the theory we shall develop is that under purely topological assumptions, $M$ has many uniquely ergodic laminations of minimal hypersurfaces.
A calibration (of codimension $1$) is a closed d1d-1-form $F$ such that $\|F\|_{L^{\infty}}=1$. For any calibration $F$ on $M$, a hypersurface $N\subset M$ is $F$-calibrated if the pullback of $F$ to $N$ is the area form on $N$. If $N$ is $F$-calibrated, then the mean curvature of $N$ is 0, and if $N$ is closed then $N$ minimizes its area in its real homology class. This brings us to our main theorem:
@@
The lamination $\lambda_{\rho}$ is the canonical lamination calibrated by $\rho$. We prove Theorem 1.2 in §4. The proof uses multiple results from [6], including the interpretation of mass-minimizing laminations in terms of functions of least gradient, functions $u$ which minimize their total variation $\int_{M}\star|\mathop{}\!\mathrm{d}u|$. We carefully note that the lamination $\lambda_{\rho}$ is not itself mass-minimizing, since it may not admit any sort of transverse measure and therefore does not have a well-defined homology class. However, any measured sublamination of $\lambda_{\rho}$ is mass-minimizing in its homology class.
To illustrate Theorem 1.2, suppose that $d=2$, so that we can identify homotopy classes of maps $M\to\mathbf{S}^{1}$ with homomorphisms $\pi_{1}(M)\to\mathbf{Z}$, which in turn can be identified with lattice points in $H^{1}(M,\mathbf{R})$. Let $\rho$ be such a lattice point; by rescaling $M$ we may assume that $\|\rho\|_{\infty}=1$. In that case, any calibration which represents $\rho$ is the derivative of a minimizing Lipschitz map in the homotopy class $\rho$, and every leaf of the canonical lamination calibrated by $\rho$ is maximally stretched by every minimizing Lipschitz map in $\rho$. Thus Theorem 1.2 is a generalization of a version of Theorem 1.1 where one works with homotopy classes of maps $M\to\mathbf{S}^{1}$ rather than $\operatorname{id}_{\Sigma_{g}}]$.
See §6.2 for a discussion of the possible generalizations of Theorem 1.2 to higher dimension and codimension.
In §5, we study measured sublaminations of canonical calibrated laminations. This is motivated by the fact that the earthquake norm, the dual of the norm induced by Thurston’s stretch metric, is not strictly convex, and its failure to be strictly convex detects the failure of geodesic laminations to be uniquely ergodic [31]. This suggests that if the stable norm is not strictly convex, then there should be canonical calibrated laminations which are not uniquely ergodic; it turns out that this is exactly what happens.
A finite Borel measure $\mu$ is transverse to a lamination $\lambda$, if $\operatorname{supp}\mu=\operatorname{supp}\lambda$ and $\mu$ is invariant under deformations which preserve the area forms of every leaf of $\lambda$. A transverse probability measure $\mu$ is ergodic if, for every Borel set $E$ which is a union of leaves of $\lambda$, either $\mu(E)=0$ or $\mu(E)=1$. The lamination $\lambda$ is uniquely ergodic, if there is a unique probability measure $\mu$ which is transverse to $\lambda$ (in which case $\mu$ must be ergodic).
@@
Auer and Bangert [3] observed that one can use a lemma of Arnoux and Levitt [2] to estimate the number of ergodic measures on sublaminations of a lamination without closed leaves. So by Corollary 1.3, the Arnoux–Levitt lemma applied to $\lambda_{\rho}$ determines the structure of $\rho^{*}$. A homology class $\alpha\in H_{d-1}(M,\mathbf{R})$ has rational direction if there exists $c>0$ such that $c\alpha$ is in the image of the map $H_{d-1}(M,\mathbf{Z})\to H_{d-1}(M,\mathbf{R})$. Let $b_{1}(M):=\dim H_{1}(M,\mathbf{Q})$ be the first Betti number.
@@
We illustrate this theorem with two examples where $\partial B$ cannot be strictly convex.
  1. (1)
    Suppose that $M$ is homeomorphic to the closed oriented surface $\Sigma_{g}$ of genus $g\geq 2$. By a theorem of Massart [38], if a maximal flat $\rho^{*}$ has a point of rational direction, then $\rho^{*}$ has at most 3g33g-3 vertices, all of which have rational direction. It follows that $\lambda_{\rho}$ consists of at most 3g33g-3 closed geodesics, plus possibly a “spiraling” part which admits no transverse measures. On the other hand, if $\rho^{*}$ has no points of rational direction, then $\rho^{*}$ has at most 2g12g-1 vertices, and $\lambda_{\rho}$ has no closed leaves.
  2. (2)
    Suppose that $d=3$, $M$ has positive scalar curvature $R\geq R_{\rm min}>0$, and $\rho^{*}$ is a maximal flat. The inradius of any stable minimal surface $N\subset M$ is $\lesssim R_{\rm min}^{-1/2}$ [34, Proposition 2.2]; this holds when $N$ is an intrinsic ball in a leaf of $\lambda_{\rho}$, since calibrated surfaces are stable minimal. So every leaf of $\lambda_{\rho}$ must be closed, and every vertex of $\rho^{*}$ must have rational direction.
@@
The analysis of $\lambda_{\rho}$ yields the following theorem on the strict convexity of the stable unit ball, which was proposed without proof by Auer and Bangert [3, Theorems 6 and 7]. The intersection product $\alpha\cdot\beta$ of two homology classes $\alpha,\beta$ is the Poincaré dual of the cup product $\operatorname{PD}(\alpha)\smile\operatorname{PD}(\beta)$, and the derived series of a group $\Gamma$ is defined by letting $\Gamma^{(0)}:=\Gamma$ and $\Gamma^{(n+1)}$ be the commutator subgroup of $\Gamma^{(n)}$.
@@
    If there is a line segment $\alpha,\beta]\subset\partial B$, then $\alpha\cdot\beta=0$.
@@
For example, suppose that $M$ has the homotopy type of a torus. Then $B$ is strictly convex and so:
  1. (1)
    $M$ has many uniquely ergodic laminations of minimal hypersurfaces. If $M=\mathbf{R}^{d}/[0,1]^{d}$, then these laminations are the irrational foliations of $M$, but what Theorem 1.6 says is that the same behavior occurs regardless of the Riemannian metric on $M$.
  2. (2)
    If $d=3$, then the discussion after Theorem 1.4 implies that $M$ does not have positive scalar curvature. This is a classic theorem of Schoen and Yau [43] which has a much more direct proof.
Since the canonical calibrated lamination $\lambda_{\rho}$ does not have to be uniquely ergodic, it is natural to ask if there is a canonical measure on $\lambda_{\rho}$. In upcoming work [18], Daskalopoulos and Uhlenbeck will show there is a favored measure on any canonical maximally stretched lamination with only closed leaves. The same proof works on $\lambda_{\rho}$, but it is quite lengthy, so we omit the proof. For each $p<\infty$, there is a unique representative $F_{p}\in L^{p}(M,\Omega^{d-1})$ of $\rho$ which is $p$-harmonic; that is,
@@
[…]
@@
Theorems 1.2 and 1.7 are explicitly based on theorems about the earthquake norm, and there are also versions of Theorem 1.4 and 1.6 for the earthquake norm proven in [31]. In §6.1 we conjecture a version of Corollary 1.3 for the earthquake norm.
@@ Acknowledgments.
This work was closely inspired by ideas in [3] of Franz Auer and Victor Bangert; I am especially grateful to Victor Bangert for allowing me to view their unpublished manuscript [4]. I also thank Georgios Daskalopoulos and Karen Uhlenbeck for helpful discussions, and James Farre, Yi Huang, Zhenhua Liu, Ben Lowe, and the anonymous referee for helpful comments on an earlier draft. This research was supported by the National Science Foundation’s Graduate Research Fellowship Program under Grant No. DGE-2040433.
@@ 2.1. Notation
Unless otherwise noted, $M$ always denotes a closed oriented Riemannian manifold of dimension $2\leq d\leq 7$. The operator $\star$ is the Hodge star on $M$. We denote the musical isomorphisms by $\sharp,\flat$. We write $H^{\ell}$ for de Rham cohomology, but never a Sobolev space, which we instead denote $W^{\ell,p}$. The second fundamental form of a submanifold $N$ is $\mathrm{I\!I}_{N}$. If $K$ is a closed compact subset of a topological vector space, $\mathcal{E}(K)$ is the set of extreme points of $K$.
@@ 2.2. Differential forms in $L^{\infty}$
Let $x\in M$, and $0\leq k\leq d-1$. Then there exists $r_{*}>0$ which depends only on $\mathrm{Riem}_{M}$ near $x$ and the injectivity radius of $x$, such that for every $0<r\leq r_{*}$ and $F\in L^{\infty}(B(x,r),\Omega^{k+1}_{\rm cl})$, there exists a Hölder continuous $k$-form AA such that $F=\mathop{}\!\mathrm{d}A$.
@@
We may choose $r_{*}$ so that the exponential map $B_{\mathbf{R}^{d}}(0,r_{*})\to B(x,r_{*})$ is a diffeomorphism which induces topological isomorphisms for every function space under consideration. Thus it is no loss to replace $B(x,r)$ with the unit euclidean ball $\mathbf{B}^{d}$. By the main theorem of [14], for every $1<p<\infty$ there is a continuous right inverse to the exterior derivative
@@
The next result is a rephrasing of [1, Theorem 1.2], and asserts that closed $L^{\infty}$ d1d-1-forms can be integrated along Lipschitz hypersurfaces.
@@
Let $\iota:N\to M$ be the inclusion of an oriented Lipschitz hypersurface. Let $\mathcal{X}$ be the space of $F\in L^{\infty}(M,\Omega^{d-1})$ such that the components of $\mathop{}\!\mathrm{d}F$ are Radon measures. Then the pullback $\iota^{*}$ of d1d-1-forms extends to a bounded linear map
@@
where the supremum ranges over all oriented $k$-dimensional submanifolds $\Sigma$. It is clear that $\|F\|_{L^{\infty}_{*}}\leq\|F\|_{L^{\infty}}$, but if $F$ is a d1d-1-form, then the converse holds as well; we shall often use this fact without comment.